# Large-time behaviour for coupled systems of Lotka-Volterra-type Fokker-Planck equations

### v1

@@ 5 Numerical tests
In this section, we provide several numerical tests to illustrate the obtained decay towards equilibrium for the system of Fokker-Planck equations (3) in terms of the energy distance (16). In particular, we illustrate the consistency of the decays through the class of structure preserving scheme introduced in [38] to approximate the evolution of the distribution functions $f_{1},f_{2}$. These methods are capable to reproduce large times statistical properties of the exact steady state with arbitrary accuracy, together with the preservation of positivity of the solution and a consistent entropy dissipation. In all the subsequent tests we initialise the parameters defined in Table 1.
@@ 5.1 Test 1: convergence to the quasi-equilibrium densities
where $m_{1}(0)=4$, $m_{2}(0)=3$. In Figure 2-3 we show the dynamics of the Energy distance for several choices of $\ell=0.6,0.7,0.8$. The evolution of the densities $f_{1},f_{2}$ have been approximated through a second order semi-implicit structure preserving scheme over the domain $[0,L]$, $L=50$ discretized with $N=1001$ gridpoints, we point the interested reader to [38] for a detailed description. The integrated the dynamics through $[0,50]$ with a time step $\Delta t=\Delta x/2$. The damping of the Energy distance can be observed, with the upper bounds derived in Section 4 indicating that the convergence becomes exponential after a transient regime.
@@
Figure 4: We represent $f_{1},f_{2}$ and $f_{1}^{q},f_{2}^{q}$ for three time steps $t_{1}=1$, $t_{2}=10$, $t_{3}=20$. The approximation of $f_{1},f_{2}$ is obtained with a 6th order semi-implicit SP scheme for the Fokker-Planck system of equations (3) under the choice of parameters in (LABEL:eq:values) and coefficients in Table 1.
@@ 5.2 Test 2: convergence to the equilibrium density for $p\in[\frac{1}{2},1]$
In Figure 5 we represent the obtained trends. The approximation of $f_{1},f_{2}$ have been obtained through a 6th order semi-implicit SP scheme for the Fokker-Planck system (3) over the domain $[0,50]$ discretized with $1001$ gridpoints and a time step $\Delta t=\Delta x/2$.
@@ Conclusion
Our analysis establishes a rigorous framework connecting kinetic descriptions of interacting populations with their macroscopic Lotka-Volterra counterparts. Through the introduction of Energy-type distances, we demonstrated the dissipative structure of the considered dynamics, which drives exponentially in time the system toward equilibrium. Thanks to the obtained bounds we quantified the rate of convergence in suitable homogeneous Sobolev spaces and we also clarified the underlying energy dissipation mechanism governing the dynamics. Overall, the proposed approach bridges kinetic modelling with Lotka-Volterra-type dynamics establishing a new perspectives in the derivation of equilibration rates in large interacting systems.

### v2

@@ 5 Numerical tests
In this section, we provide several numerical tests to illustrate the obtained decay towards equilibrium for the system of Fokker-Planck equations (3) in terms of the energy distance (16). In particular, we illustrate the consistency of the decays through the class of structure preserving scheme introduced in [39] to approximate the evolution of the distribution functions $f_{1},f_{2}$. These methods are capable to reproduce large times statistical properties of the exact steady state with arbitrary accuracy, together with the preservation of positivity of the solution and a consistent entropy dissipation. In all the subsequent tests we initialise the parameters defined in Table 1. In all the subsequent tests we considered a domain $[0,L]$, $L=50$, discretized by means of $N=1001$ gridpoints. The discretization of the considered time interval has time step $\Delta t>0$ such that $\Delta t=\Delta x/2$ to comply with positivity preservation os semi-implicit schemes discussed in [39]. At the boundary $x=0$ we always considered a no-flux boundary condition
@@ 5.1 Test 1: convergence to the quasi-equilibrium densities
where $m_{1}(0)=4$, $m_{2}(0)=3$. In Figure 2-3 we show the dynamics of the Energy distance for several choices of $\ell=0.6,0.7,0.8$. The evolution of the densities $f_{1},f_{2}$ have been approximated through a second order semi-implicit structure preserving scheme over the domain $[0,L]$, $L=50$ discretized with $N=1001$ gridpoints, we point the interested reader to [39] for a detailed description. We integrated the dynamics over the time horizon $[0,50]$ with a time step $\Delta t=\Delta x/2$. The damping of the Energy distance can be observed, with the upper bounds derived in Section 4 indicating that the convergence becomes exponential after a transient regime.
@@
Figure 4: We represent $f_{1},f_{2}$ and $f_{1}^{q},f_{2}^{q}$ for three time steps $t_{1}=1$, $t_{2}=10$, $t_{3}=20$. The approximation of $f_{1},f_{2}$ is obtained with a 2nd order semi-implicit SP scheme for the Fokker-Planck system of equations (3) under the choice of parameters in (LABEL:eq:values) and coefficients in Table 1.
@@ 5.2 Test 2: convergence to the equilibrium density for $p\in[\frac{1}{2},1]$
In Figure 5 we represent the obtained trends. The approximation of $f_{1},f_{2}$ have been obtained through a 6th order semi-implicit SP scheme for the Fokker-Planck system (3) over the domain $[0,50]$ discretized with $1001$ gridpoints and a time step $\Delta t=\Delta x/2$. We observe that the computed rate is suboptimal and apparently the system of equations converge faster than the obtained exponential rate.
@@ Conclusion
Our analysis establishes a rigorous framework connecting kinetic descriptions of interacting populations with their macroscopic Lotka-Volterra counterparts. Through the introduction of Energy-type distances, we demonstrated the dissipative structure of the considered dynamics, which drives exponentially in time the system toward equilibrium. Thanks to the obtained bounds we quantified the rate of convergence in suitable homogeneous Sobolev spaces and we also clarified the underlying energy dissipation mechanism governing the dynamics. Overall, the proposed approach bridges kinetic modelling with Lotka-Volterra-type dynamics establishing a new perspectives in the derivation of equilibration rates in large interacting systems. More generally, we expect that the developed techniques can be fruitfully applied to a vast class of modelling settings.
Data Accessibility
All the sources and data of the manuscript have been provided in the following repository https://doi.org/10.5281/zenodo.18756863